### v1

# RLGNet: Repeating-Local-Global History Network for Temporal Knowledge Graph Reasoning

Abstract
Temporal Knowledge Graph (TKG) reasoning is based on historical information to predict the future. Therefore, parsing and mining historical information is key to predicting the future. Most existing methods fail to concurrently address and comprehend historical information from both global and local perspectives. Neglecting the global view might result in overlooking macroscopic trends and patterns, while ignoring the local view can lead to missing critical detailed information. Additionally, some methods do not focus on learning from high-frequency repeating events, which means they may not fully grasp frequently occurring historical events. To this end, we propose the Repetitive-Local-Global History Network(RLGNet). We utilize a global history encoder to capture the overarching nature of historical information. Subsequently, the local history encoder provides information related to the query timestamp. Finally, we employ the repeating history encoder to identify and learn from frequently occurring historical events. In the evaluation on six benchmark datasets, our approach generally outperforms existing TKG reasoning models in multi-step and single-step reasoning tasks.
1 Introduction
TKG is a structured yet highly complex knowledge system. In TKG, each fact is represented by a quadruple, including subject, relation, object, and timestamp, thereby providing detailed and accurate temporal information. TKG allow us to understand the relationships between entities and capture the changes in these relationships via timestamps.
TKG extrapolation aims to predict new facts at a future timestep based on the ground truth facts within the known timeframe. Extrapolation is further classified into two types: (1) Single-step reasoning predicts future events using known ground truth facts from past timestamps within the test period. (2) Multi-step reasoning forecasts without relying on ground truth facts from the test period. By analyzing past events and identifying factors that influence event progression, we can infer the probabilities of emergent future facts. However, since future events contain unknown factors, TKG extrapolation is a challenging problem(?).
Figure 1: A figure to explain repeating, local, and global facts. ”Frequency” represents the number of occurrences of that type of event, ”Interval” represents the time difference from the last occurrence of that type of event to the time of query. The red arrows represent the facts most relevant to the query.
Predicting the future is a task that requires a deep analysis of historical data, the key to which is to discern possible patterns and trends from it. In previous work, it’s generally believed that the facts closer to the query time are more important. However, this method may cause us to overlook some facts that have the potential to impact future trends, because these facts are not necessarily always close to the query time. As shown in Figure 1, (United States, Negotiations, ?) does not appear in the local history. According to historic recurrence(?), historical facts could have repeated or even cyclical patterns. As illustrated in the same Figure 1, (United States, Negotiations, North Korea) has occurred 95 times. Therefore, we need to consider the repetitiveness, locality, and globality of historical information. Moreover, some work has shown that not every fact in the TKG is crucially important for prediction. To improve the accuracy of prediction, we need to selectively pay attention to the facts that are most relevant to the query, as the fact indicated by the red arrow in Figure 1. Therefore, finding a method that is based on the query and simultaneously considers repeating, local and global historical information is of paramount importance for extrapolation in TKG.
To address this challenge, we design and propose a model called Repeating-Local-Global History Network (RLGNet). In this model, we design repeating, local, and global history encoders respectively based on the query-oriented method. In the local history encoder, the model views the TKG as a Knowledge Graph (KG) sequence and models the facts in adjacent timestamps to obtain local historical information. In the global history encoder, the model aggregates all facts using an attention mechanism to gather global historical information. Finally, building on the foundation of local and global historical encoders, we use a repeating history encoder to further increase the scores of repetitive events. Overall, this paper makes the following contributions:
  1. 1.
    We designed three different levels of encoders to deal with the repetitive, local, and global historical information respectively.
  2. 2.
    We explored the contributions of repeating, global and local historical information to predictive performance in two different extrapolation settings.
  3. 3.
    We conducted extensive experiments on six public TKG datasets, proving the model’s effectiveness in entity prediction.
2 Related Work
Static KG Reasoning: Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities. Another category is semantic-matching-based models, like DistMult(?) and RESCAL(?). There are also models based on Convolutional Neural Networks (CNNs), such as ConvE(?) and Conv-TransE(?), which represent entities and relations using matrices processed by convolutional kernels. Similarly, Graph Convolutional Network (GCN) models, including R-GCN(?) and VR-GCN(?), stand out for their capacity to integrate graph structures with node features. However, these models focus on static KG, and their predictive capabilities for future events are limited.
Temporal KG Reasoning: TKG reasoning has two settings: extrapolation and interpolation. The interpolation setting aims at predicting missing historical facts rather than future events. In contrast, this paper focuses on reasoning within the extrapolation setting to forecast future facts using historical data. Know-Evolve(?) and DyREP(?) model the occurrence of facts in TKG using a temporal point process. Glean(?) enriches factual features by utilizing unstructured text information. CyGNet(?) captures and understands historical trends and patterns through a replicating generative mechanism. CENET(?) distinguishes events as historical and non-historical, using this classification for contrastive learning. Some models, such as RE-GCN(?), utilize Graph Convolutional Networks (GCNs) to simulate the evolutionary process of knowledge graphs, thereby capturing and learning the dynamic properties of entities and relations. TiRGN(?) uses a local-global historical approach for reasoning. CEN(?) and CluSTeR(?) utilize reinforcement learning to capture and learn the connections between entities and relationships in the knowledge graph. TLogic(?) utilizes temporal logic rules to constrain the predicted paths for queries. TANGO(?) utilizes neural ordinary differential equations to model the structural information of each entity. However, these models often focus on partial aspects of historical information without combining query data with historical details. Consequently, they can’t accurately capture recurring, local, and global facts.
3 Preliminaries
TKG represent events temporally as snapshots. Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}\}$. Each fact in the snapshot is a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{T}$. It represents a relationship $r$ between subject entity $s$ and object entity $o$ at time $t$. For each quadruple $(s,r,o,t)$,an inverse relation quadruple$(o,r^{-1},s,t)$ is often added to the dataset. The entity prediction task aims to predict the missing entity in a query $q=(s_{q},r_{q},?,t_{q})$. We denote all entity embeddings in dimension $d$ with $H\in\mathbb{R}^{|\mathcal{E}|\times d}$, and relation embeddings similarly with $R\in\mathbb{R}^{|\mathcal{R}|\times d}$. In each encoder, the initial weights of H and R are randomly generated. The specific entity and relation are represented as $h\in\mathbb{R}^{d}$and $r\in\mathbb{R}^{d},$ respectively.
Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},o_{q},o,t)\in\mathcal{G}_{t}\}$. The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range.
$\mathcal{C}_{q}^{t_{1}:t_{2}}=\bigcup_{i=t_{1}}^{t_{2}}\mathcal{C}_{q}^{i}$ (1)
Since not all candidate entities greatly affect the prediction, we only keep the top $k$ most frequent entities, denoted as $C_{q,top_{k}}^{1:t_{q}-1}$. Candidates in the set usually represent repeating facts, we construct $\mathcal{G}_{rep}$, containing all tuples in the period where o is a top $k$ candidate before $t-1$. $\mathcal{G}_{rep}$ can be formalized as:
$\mathcal{G}_{rep}=\{q=(s_{q},r_{q},o_{q},t_{q})\in\mathcal{G}_{raw}|o\in\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}\}$ (2)
$\mathcal{G}_{rep}$ essentially extracts prominent, repetitive events from historical data.
Figure 2: The overall architecture of RLGNet. The upper half is the local history encoder which learns local historical information through KG sequences of adjacent timestamps. The lower half comprises the repeating and global history encoders, learning repetitive and global historical information respectively by statistically querying candidate entities. The content within the dashed box represents the entity most relevant to the query. The left side sums the scores of the three parts to get the final prediction score.
4 Model Overview
The overall framework of the model is illustrated in the Figure 2. This model can be divided into three core submodules: the local history encoder, the global history encoder and the repeating history encoder. The local history encoder captures local facts by exploring the structural features and historical dependencies the sequence of KG at adjacent timestamps. The global history encoder captures global facts by examining the relevant facts across all previous timestamps and capturing entities and relationships that may not be evident at adjacent timestamps. Furthermore, these two encoders use a scoring decoder to evaluate the facts. The primary task of the repeating history encoder is to enhance its ability to predict repetitive facts by encoding frequently occurring historical events. All encoders utilize time vector to encode time and frequency. By utilizing these three encoders, the model is capable of making accurate predictions in response to queries.
4.1 Scoring Decoder
Inspired by the work of ConvE, we designed this scoring decoder and used it in the next two decoders. First, the model reshapes a one-dimensional input into a two-dimensional matrix. Then, it processes this matrix with a CNN. After that, it flattens the resulting data and outputs the embeddings for the target entity through a fully connected layer.
$\varphi(H,V_{1}\dots V_{k})=\sigma(ConvE(V_{1},\dots,V_{k}))\cdot H$ (3)
$\sigma$ represents the ReLu function. $V_{i}(1\leq i\leq k)$are input vectors, $\varphi(H,V_{1},\dots,V_{k})\in\mathbb{R}^{|\mathcal{E}|}$ represents predicted entity scores.
4.2 Time Vector
In terms of time, the module employs two types of temporal encoding vectors, periodic(?) and aperiodic, to represent time.
$\textbf{v}_{1}(t)&=\cos(W_{1}^{t}t+b_{1})\\ \textbf{v}_{2}(t)&=\tanh(W_{2}^{t}t+b_{2})\end{aligned}\right.$ (4)
Then we concatenate the two vectors to obtain the final vector representation.
$\textbf{V}(t)=[\textbf{v}_{1}(t);\textbf{v}_{2}(t)]$ (5)
$W_{1}^{t},W_{2}^{t},b_{1},b_{2}$ are learnable parameters and the vectors $\textbf{v}_{1}(t)$ and $\textbf{v}_{2}(t)$ have the same dimensionality, [;] denotes the vector concatenation operation, $\textbf{V}(t)$ is the time vector.
4.3 Local History Encoder
This module captures local facts by focusing on adjacent history. For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}_{t_{q}-m},\dots,\mathcal{G}_{t_{q}-1}\}$ related to this query to obtain the structural features of the subgraph. For this purpose, we employ GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU)(?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities. Firstly, we use GCN with $\omega$ layers to get the entity representations at the current moment.
$h_{o}^{t,l+1}=\sum\limits_{(s,r,o)\in\mathcal{G}_{t}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{i}^{t,l},r_{t})+W_{2}^{l}h_{o}^{l}$ (6)
$\phi(h_{i}^{t,l},r_{t})=[h_{i}^{t,l};r_{t};h_{i}^{t,l}+r_{t};h_{i}^{t,l}\cdot r_{t}]$ (7)
The symbol [$\cdot$] represents the Hadamard product. $N_{o}^{t}$ represents neighbors of node $o$ at time $t$. The embedded representation of node $o$ at the $l$-th layer is denoted as $h_{o}^{t,l}\in\mathbb{R}^{d}$. The aggregation and self-loop parameters of the $l$-th layer are shown as $W_{1}^{l}$ and $W_{2}^{l}$. The edge relation type is presented as $r_{t}$. Specifically, when $\omega=0$, i.e., $h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$. Entity representations are combined with their temporal representations and then predicted for the next moment with GRU:
$H_{t+1}=GRU([H_{t};\textbf{V}_{1}(t_{q}-t)],H^{{}^{\prime}})$ (8)
$\textbf{V}_{1}(t_{q}-t)\in\mathbb{R}^{d_{2}}$ is the time vector. $H^{{}^{\prime}}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the final GCN layer output. The attention mechanism aggregates candidate entities over a history of length $m$. Node aggregation for query $q$ is:
$C_{q}^{t}=\sum\limits_{i\in C_{q}^{t}}\frac{1}{|C_{q}^{t}|}h_{i}^{{}^{\prime}}$ (9)
$C_{q}^{t}$ represents query $q$’s candidate entity representation at time $t$. Next, the module calculates attention weights $a_{q}^{t}$ for the query at time $t$:
$a_{q}^{t}=W_{2}\sigma(W_{1}[h_{q}^{{}^{\prime}};r_{q};\textbf{V}_{1}(t_{q}-t);C_{q}^{t}])$ (10)
$W_{1}$, and $W_{2}$ are learnable parameters. After calculating attention weights, these are used to aggregate candidate entities for query q:
$C_{q}^{{}^{\prime}}=\sum_{i=1}^{m}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=1}^{m}\exp{(a_{q}^{j})}}$ (11)
The candidate entity representation $C_{q}^{{}^{\prime}}$ for query $q$. The module then calculates the entity score for query $q$’s local history encoder:
$score_{loc}^{q}=\varphi_{loc}(H^{{}^{\prime}},h_{q}^{{}^{\prime}},r_{q},C_{q}^{{}^{\prime}})$ (12)
Eq (12) can be viewed as an instance of Eq (3).
4.4 Global History Encoder
The global history encoder is designed to consider and extract information from a global perspective in order to capture global facts. For each query, we count the frequencies and the recent appearances of candidate entities. The module then uses an attention mechanism to integrate this information about the candidate entities. First, the attention scores for the time-related candidate entities of the current query are calculated:
$\mathcal{A}_{1}(q,i)=\sigma(W_{3}[h_{q};r_{q}])^{T}\sigma(W_{4}[\textbf{V}_{2}(t_{q}-t_{q}^{i});h_{i}])$ (13)
The attention scores for frequency-related candidates is similarly:
$\mathcal{A}_{2}(q,i)=\sigma(W_{5}[h_{q};r_{q}])^{T}\sigma(W_{6}[\textbf{V}_{3}(cnt^{i}_{q});h_{i}])$ (14)
Considering the similarities between the characteristics of frequency and time, we employ a time vector to represent frequency. $W_{4}\in\mathbb{R}^{d\times 2d},W_{3}\in\mathbb{R}^{d\times 2d},W_{5}\in\mathbb{R}^{d\times 2d}$, and $W_{6}\in\mathbb{R}^{d\times 2d}$ are learnable parameters. $t_{q}^{i}$ and $cnt_{q}^{i}$ represent the frequency and last occurrence of candidate entity $i$ for query q. Additionally, taking computational resources into consideration, we only selected $top_{k}^{all}$ candidate entities. Following, we normalize their respective scores and multiply by the candidate entity representations to obtain representations $C_{q}^{gap}$ and $C_{q}^{cnt}$ for time-relevant and frequency-relevant candidates, respectively.
$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (15)
$C_{q}^{cnt}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{2}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{2}(q,j))}$ (16)
The module then calculates the entity score for the global history encoder:
$score_{glo}^{q}=\varphi_{glo}(H,h_{q},r_{q},\textbf{V}_{4}(t_{q});C_{q}^{gap};C_{q}^{cnt})$ (17)
$h_{q}$ and $r_{q}$ represent the query entity and relation, Eq (17) can be viewed as an instance of Eq (3).
4.5 Repeating History Encoder
The Repeating History Encoder is designed to improve prediction by increasing the weight of historical events. We first filter out and retain all the repetitive facts. Then, only the candidate entities of facts that rank in the top($top_{k}$) in terms of frequency are retained. In this module, the frequency of the candidate entities is also represented using a time vector. The repeating history encoder’s score function is a multilayer perceptron (MLP):
$\textbf{MLP}(x)=W_{7}\cdot\sigma(W_{8}\cdot\sigma(W_{9}\cdot x))$ (18)
$W_{7}\in\mathbb{R}^{1\times d}$, $W_{8}\in\mathbb{R}^{d\times 3d}$, and $W_{9}\in\mathbb{R}^{3d\times 5d}$ are learnable parameters. The extra score for candidates is calculated as follows:
$score_{rep}^{q,i}=\textbf{MLP}([h_{q};r_{q};\textbf{V}_{5}(t_{q});h_{i};\textbf{V}_{6}(cnt_{q}^{i})])$ (19)
It should be noted that when entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$ the $score_{rep}^{q,i}$ is 0.
4.6 Loss Function
The model uses the cross-entropy loss function as a standard to measure loss, with the following formula:
$\mathcal{L}(score_{i},\mathcal{G}_{\mathcal{F}(i)})=\sum_{(s,r,o,t)\in\mathcal{G}_{\mathcal{F}(i)}}y_{t}\log P_{i}(o|s,r,t)$ (20)
where $P_{i}(o|s,r,t)=softmax(score_{i})$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$is the label vector, where an element is 1 if the fact occurs, or 0 otherwise. Since each encoder in the model is trained separately, each one has a different loss function. When $i=rep$ , $\mathcal{F}(i)$ equals $rep$. And when $i\in\{loc,glo\}$, $\mathcal{F}(i)$ retrieves $raw$.
4.7 Final Score
The final model score combines scores from the repeating, local and global history encoders.
$\begin{split}score_{fin}^{q}=&\ \alpha\cdot score_{loc}^{q}+(1-\alpha)\cdot score_{glo}^{q}\\ &+score_{rep}^{q}\end{split}$ (21)
$\alpha\in[0,1]$ is a hyperparameter.
Model ICEWS18 ICEWS14 ICEWS05-15
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
Single-Step xERTE 29.31 21.03 33.51 46.48 40.79 32.70 45.67 57.30 46.62 37.84 52.31 63.92
RE-GCN 32.62 22.39 36.79 52.68 42.00 31.63 47.20 61.65 48.03 37.33 53.90 68.51
TITer 29.98 22.05 33.46 44.83 41.73 32.74 46.46 58.44 47.60 38.29 52.74 64.86
TiRGN 33.66 23.19 37.99 54.22 44.04 33.83 48.95 63.84 50.04 39.25 56.13 70.71
CEN 31.50 21.70 35.44 50.59 42.20 32.08 47.46 61.31 46.84 36.38 52.45 67.01
RETIA 32.43 22.23 36.48 52.94 42.76 32.28 47.77 62.75 47.26 36.64 52.90 67.76
RLGNet 34.96 24.68 39.22 55.09 46.15 36.16 51.17 65.12 50.56 40.34 56.05 70.18
Multi-Step CyGNet 26.07 16.76 29.54 44.43 34.80 25.34 39.05 53.09 38.17 27.93 43.01 57.89
RE-GCN 28.44 19.03 31.96 46.86 37.68 28.00 41.81 56.87 38.74 28.50 43.60 58.52
TiRGN 28.85 19.18 32.58 47.78 38.37 28.80 42.50 56.94 39.97 29.44 44.76 60.92
CENET 27.40 18.91 30.26 44.36 35.62 27.10 38.81 52.31 39.92 30.21 44.14 59.09
RLGNet 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12 40.83 30.06 45.91 61.93
Table 1: Performance (in percentage) on on ICEWS18, ICESW14, and ICEWS05-15.
Model WIKI YAGO GDELT
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
Single-Step xERTE 73.60 69.05 78.03 79.73 84.19 80.09 88.02 89.78 19.45 11.92 20.84 34.18
RE-GCN 78.53 74.50 81.59 84.70 82.30 78.83 84.27 88.58 19.69 12.46 20.93 33.81
TITer 73.91 71.70 75.41 76.96 87.47 84.89 89.96 90.27 18.19 11.52 19.20 31.00
TiRGN 81.65 77.77 85.12 87.08 87.95 84.34 91.37 92.92 21.67 13.63 23.27 37.60
CEN 78.93 75.05 81.90 84.90 83.49 79.77 85.85 89.92 20.39 12.96 21.77 34.97
RETIA 78.59 74.85 81.39 84.58 81.04 77.00 83.31 88.62 20.12 12.76 21.45 34.49
RLGNet 82.43 78.86 85.65 87.17 89.69 87.05 92.15 93.00 25.09 16.95 27.42 40.87
Multi-Step CyGNet 58.44 53.03 62.24 67.46 68.60 60.97 73.58 83.16 19.11 11.90 20.31 33.12
RE-GCN 62.05 58.95 63.89 67.39 70.05 65.76 72.70 77.16 19.62 12.47 20.86 33.48
TiRGN 64.04 60.72 66.52 68.96 78.51 74.01 82.74 84.76 19.87 12.46 21.21 34.25
CENET 57.52 51.99 61.93 66.29 69.90 64.01 73.04 82.65 - - - -
RLGNet 64.34 61.03 66.71 69.51 80.17 76.52 83.57 84.96 20.81 13.34 22.32 35.38
Table 2: Performance (in percentage) on WIKI, YAGO, and GDELT.
5 Experiments
5.1 Setup
Dataset $||\mathcal{E}|$$|\mathcal{R}|$ Train Valid Test Time gap
ICE18 23,033 256 373,018 45,995 49,545 24 hours
ICE14 7,128 230 63,685 13,823 13,222 24 hours
ICE15 10,488 251 368,868 46,302 46,159 24 hours
WIKI 12,554 24 539,286 67,538 63,110 1 year
YAGO 10,623 10 161,540 19,523 20,026 1 year
GDELT 7,691 240 1,734,399 238,765 305,241 15 mins
Table 3: Statistics of the datasets.
Datasets. We use six TKG datasets to evaluate the model’s effectiveness in entity prediction task, including ICEWS14(?), ICEWS18(?), and ICEWS05-15(?) from the Integrated Crisis Early Warning System (ICEWS) and the event-driven GDELT(?) dataset. Public datasets WIKI(?) and YAGO(?) are also included. All datasets are time-partitioned into Training (80%), Validation (10%), and Test (10%). More details on datasets are in Table 3.
Evaluation Metrics. To assess TKG reasoning performance, we used Mean Reciprocal Rank (MRR) and Hits@k metrics. MRR calculates average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings. Previous studies have indicated that traditional filtering settings are flawed(?). Thus, we report the experimental results after using the time-aware filtering settings.
Baselines. RLGNet is compared with six baseline models on single-step reasoning task, including xERTE(?), RE-GCN(?), TITer(?), TiRGN(?), CEN(?), and RETIA(?). Since some models are not designed for multi-step reasoning task, we have selected a few models to report their performance on multi-step reasoning task, including CyGNet(?), RE-GCN(?), TiRGN(?), and CENET(?).
Implementation Details. For all datasets, we set the vector dimension $d$ to 200, and the time vector dimension $d_{2}$ in the local history encoder to 48. The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200. The number of GCN layers $\omega$ is set to 1. For ICEWS18, ICEWS14, ICEWS05-15, GDELT, WIKI, and YAGO, the adjacent history length $m$ is set to 10, 10, 15, 10, 1, and 1, respectively. The hyperparameter $\alpha$ is set to 0.8 in ICEWS, 0.9 in YAGO and WIKI, and 0.1 in GDELT. At the same time, a static graph constraint similar to that of RE-GCN is also added to ICEWS. Adam is used for parameter learning, with a learning rate set to 0.001. In the local history encoder, we employ StepLR to adjust the learning rate. In this setup, the value of gamma is set to 0.8. For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, step-size is 10. Please note that each encoder is trained independently, which results in them not sharing weights.
5.2 Results
The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases. On six benchmark datasets, especially on the GDELT and ICEWS14 datasets, the MRR scores of RLGNet have increased by 3.42% and 2.11%, respectively. The primary reason for this improvement is that RLGNet not only considers repeating, local, and global historical information, but also integrates query information, which is not simultaneously considered by other models.
In single-step reasoning performance comparison, RLGNet demonstrates stronger performance in most cases. In the TKG extrapolation task, methods that consider query information usually achieve higher H@1 results, in comparison to the methods that ignore it(?). However, these methods may perform slightly worse on other metrics like H@3 and H@10. This could be because methods that don’t consider query information may introduce answers that don’t fully match the specific query, leading to inferior performance on H@1. This is also why, in the single-step reasoning task on the ICEWS05-15 dataset, even though TiRGN surpasses RLGNet by 0.08% in H@3 and 0.53% in H@10, RLGNet still outperforms TiRGN by 1.09% in H@1.
In multi-step reasoning task, the model is unable to know the ground truth facts during the test period. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. This is why global historical information becomes more important in multi-step reasoning task compared to single-step reasoning. TiRGN and RE-GCN has not fully considered global historical information, while CyGNet and CENET have overlooked local historical information. In contrast, RLGNet effectively integrates repetitive, local, and global historical information based on query information. Therefore, it achieves the best performance in multi-step reasoning task.
5.3 Ablation Study
Model ICEWS18 ICWS14 ICEWS05-15 WIKI YAGO GDELT
Single Multi Single Multi Single Multi Single Multi Single Multi Single Multi
Glo 32.54 29.10 42.86 38.51 45.41 39.05 59.46 52.94 71.85 60.92 24.72 20.63
Loc 34.08 28.19 44.68 37.09 50.04 38.88 81.72 63.86 87.22 77.84 21.57 19.42
Glo + Loc 34.52 29.13 45.56 38.42 50.23 39.97 81.75 63.88 87.38 78.39 24.73 20.72
Glo + Rep 33.19 29.56 43.74 38.64 46.57 40.14 66.75 56.98 68.85 59.29 25.11 20.74
Loc + Rep 34.93 29.39 45.91 38.41 50.75 40.21 82.35 64.33 89.65 79.50 23.27 20.41
RLGNet 34.96 29.90 46.15 39.06 50.56 40.83 82.43 64.34 89.69 80.17 25.09 20.81
Table 4: The MRR (in percentage) results of the ablation studies.
$top_{k}$ ICE18 ICE14 ICE15 WIKI YAGO GDELT 5 34.49 51.94 42.77 26.69 84.05 77.67 10 40.29 58.26 47.10 34.26 90.76 85.37 20 44.63 62.90 49.82 42.35 92.63 86.96 30 46.58 64.80 50.69 47.05 92.73 87.03 100 49.76 67.59 51.75 58.41 92.73 87.04 $\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
Table 5: The proportion of repeating facts when $top_{k}$ is set to different values.
To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks on six benchmark datasets. The results are shown in Tables 4. Our ablation studies indicated that the local history encoder (Loc) and the global history encoder (Glo) do not always improve results. Specifically, the use of Glo and Loc modules led to a decrease in performance in single-step reasoning tasks in the ICEWS05-15 and GDELT datasets. However, considering that their improvement in other tasks far outweighs the decline in these two subtasks, we thus believe Glo and Loc are beneficial. On the other hand, the Repetitive History Encoder (Rep) can improve the final performance across various tasks, suggesting that events usually exhibit a degree of repetition. As shown in Table 5, the proportion of repetitive events in the dataset well corroborates this point. Overall, while Loc and Glo may not perform ideally in some specific tasks, these three historical information encoding methods can all contribute to better prediction results.
5.4 Hyperparameter Analysis
Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
Effect of $\alpha$ Values in RLGNet. When processing multi-step and single-step reasoning tasks, global and local encoders contribute differently to the answers, as shown in Figure 3. In single-step reasoning task, we find that on the ICEWS dataset, the performance is best when $\alpha$ is between 0.6 and 0.8. On the WIKI and YAGO datasets, the performance peaks when $\alpha$ is between 0.8 and 1. On the GDELT dataset, the optimum range for $\alpha$ is between 0 and 0.2. Therefore, we set $\alpha$ to 0.8 and 0.1 in ICEWS and GDELT respectively, and to 0.9 in both WIKI and YAGO. In multi-step reasoning task, in order to achieve optimal performance, the value of $\alpha$ is usually smaller than the value of $\alpha$ in single-step inference tasks. This is because the local history encoder can more easily capture the sequential pattern of facts when the baseline truth is known. It’s noteworthy that in the GDELT dataset, the contribution of global historical information always exceeds that of local historical information. This is due to the GDELT dataset containing more abstract entity concepts, making it more difficult for the model to capture the sequential patterns of facts(?).
Figure 4: The impact of $\omega$ and $top_{k}$ on MRR(in percentage) results in ICEWS14 and YAGO.
Effect of $\omega$ and $tok_{k}$ Values in RLGNet. Figure 4 shows the impact of $\omega$ and $top_{k}$ on the MRR results in the ICEWS14 and YAGO datasets. As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines, so we set $\omega$ to 1 in all datasets. In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset. Therefore, we set $top_{k}$ for all datasets to 20.
6 Conclusion
In this paper, we propose RLGNet for TKG reasoning. RLGNet captures the characteristics of corresponding facts by using repeating, local, and global historical encoders respectively. In the future, finding a superior strategy to balance the contributions of global and local historical information is a worthwhile research problem. Experiments on six benchmark datasets prove that RLGNet outperforms existing models in most extrapolation tasks.
7 Limitations
In this paper, we have proposed a TKG prediction model based on three encoders. However, our approach has certain limitations:
  • 1.
    Due to the constraints of computational resources, the three encoders need to train independently. This method prevents them from sharing weights among themselves, potentially affecting the predictive performance of the model.
  • 2.
    In multi-step and single-step reasoning tasks, the contributions of global and local historical information should indeed be different. We have adopted a weighted average strategy to alleviate this issue to some extent, but choosing a specific hyperparameter $\alpha$ cannot achieve optimal performance on both multi-step and single-step reasoning tasks at the same time.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.61866008).